### v2

# Do Language Models Understand Time?

Abstract.
Large language models (LLMs) have revolutionized video-based computer vision applications, including action recognition, anomaly detection, and video summarization. Videos inherently pose unique challenges, combining spatial complexity with temporal dynamics that are absent in static images or textual data. Current approaches to video understanding with LLMs often rely on pretrained video encoders to extract spatiotemporal features and text encoders to capture semantic meaning. These representations are integrated within LLM frameworks, enabling multimodal reasoning across diverse video tasks. However, the critical question persists: Can LLMs truly understand the concept of time, and how effectively can they reason about temporal relationships in videos? This work critically examines the role of LLMs in video processing, with a specific focus on their temporal reasoning capabilities. We identify key limitations in the interaction between LLMs and pretrained encoders, revealing gaps in their ability to model long-term dependencies and abstract temporal concepts such as causality and event progression. Furthermore, we analyze challenges posed by existing video datasets, including biases, lack of temporal annotations, and domain-specific limitations that constrain the temporal understanding of LLMs. To address these gaps, we explore promising future directions, including the co-evolution of LLMs and encoders, the development of enriched datasets with explicit temporal labels, and innovative architectures for integrating spatial, temporal, and semantic reasoning. By addressing these challenges, we aim to advance the temporal comprehension of LLMs, unlocking their full potential in video analysis and beyond. Our paper’s GitHub repository can be found here.
Language language models, Videos, Temporal, Interaction
journalyear: 2025copyright: ccconference: Companion Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Companion Proceedings of the ACM Web Conference 2025 (WWW Companion ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3701716.3717744isbn: 979-8-4007-1331-6/2025/04ccs: Computing methodologies Motion captureccs: Computing methodologies Neural networksccs: Information systems Language modelsccs: Information systems World Wide Web
Figure 1. Do language models understand time? In the kitchen arena, where burritos are rolled, rice waits patiently, and sauce steals the spotlight, LLMs try their best to keep up. Captions flow like a recipe—precise and tempting—but can they truly tell the difference between prepping, cooking, and eating? After all, in cooking, timing isn’t just everything—it’s the secret sauce!
1. Introduction
Large language models (LLMs) have brought transformative advancements to artificial intelligence (AI), excelling across a wide array of tasks in natural language processing and computer vision (Bugliarello et al., 2021; Ghosh et al., 2024; Xu and Poo, 2023). Their ability to understand and generate human-like language has enabled groundbreaking applications, from machine translation to image and video captioning (Kim et al., 2024) (see Figure 1, frames from EPIC-KITCHENS-100 (Damen et al., 2022)). More recently, the integration of LLMs into video processing has sparked significant interest, leading to advances in tasks such as action recognition (Wang et al., 2022b, 2024h), anomaly detection (Wang et al., 2019b; Zhang et al., 2024c; Zhu et al., 2024), and video summarization (Islam et al., 2024; Yang et al., 2024a; Liu et al., 2024d; Zohar et al., 2024). However, videos pose unique challenges compared to other modalities due to their dual reliance on both spatial and temporal information (Chen et al., 2024d). Unlike static images, videos capture the dimension of time, embedding sequential dynamics that demand sophisticated reasoning (Vondrick et al., 2016; Liu et al., 2018). Similarly, unlike textual data, videos involve rich, complex visual elements that require intricate modeling (Wang et al., [n. d.]; Chen et al., [n. d.]).
Despite these advancements, a fundamental question remains unresolved: Do language models truly understand the concept of time? Temporal reasoning, the ability to comprehend and infer relationships between events over time, is essential for video-based tasks such as causal inference (Chen et al., 2021; Lin et al., 2022), event prediction (Huang et al., 2024; Guo et al., 2024b), and understanding action progression (Wang, 2023). While pretrained video encoders provide LLMs with spatiotemporal embeddings and text encoders contribute semantic insights, the fusion of these components often lacks the nuanced understanding of time required for more advanced applications (Ding et al., 2024). Current methods rely heavily on pretrained encoders and dataset-specific tuning (Qin et al., 2022; Wang et al., [n. d.]; Chen et al., [n. d.], 2024d), raising questions about the generalizability and scalability of these approaches. Table 1 provides a summary of recent video-LLMs, detailing the visual encoders they use and their mechanisms for interacting with these encoders.
This work aims to address these gaps and critically examines the role of LLMs in understanding temporal dynamics in video data. We focus on the interplay between visual (image/video) encoders and LLMs, exploring how effectively they bridge the gap between raw spatiotemporal features and high-level temporal reasoning. By tackling these issues, we seek to shed light on the limitations of existing approaches and inspire innovations in LLM-based video understanding. Our contributions are threefold:
  1. i.
    We provide a detailed review of LLM applications in video processing, with a particular focus on their ability to comprehend temporal concepts, highlighting the state of the art and identifying key limitations.
  2. ii.
    We analyze the shortcomings of existing LLM-based approaches, particularly their reliance on pretrained encoders and the challenges posed by video datasets, such as lack of temporal annotations and biases towards short-term dependencies.
  3. iii.
    We propose actionable pathways for advancing LLMs’ temporal understanding, emphasizing joint training strategies, better dataset design, and improved alignment between spatiotemporal and semantic features.
Model Venue Visual encoder Other modality encoders Interaction / Fusion mechanism Description
Flamingo (Alayrac et al., 2022) NeurIPS 2022 Normalizer-Free ResNet(Brock et al., 2021) Text: Chinchilla(Hoffmann et al., 2022) Perceiver Resampler & Gated XATTN-DENSE Visual-language model.
LaViLa (Zhao et al., 2023) CVPR 2022 TimeSformer(Bertasius et al., 2021) Text: 12-layer Transformer Cross-attention modules Large-scale language model.
mPLUG-2 (Xu et al., 2023) ICML 2023 CLIP ViT-L/14(Radford et al., 2021) Text: BERT(Devlin et al., 2019) Universal layers & cross-attention modules Modularized multi-modal foundation model.
Vid2Seq (Yang et al., 2023) CVPR 2023 CLIP ViT-L/14(Radford et al., 2021) Text: T5-Base(Raffel et al., 2020) Cross-modal attention Sequence-to-sequence video-language model.
Video-LLaMA (Zhang et al., 2023a) EMNLP 2023 CLIP ViT-G(Radford et al., 2021) Text: Vicuna(Chiang et al., 2023), Audio: ImageBind(Girdhar et al., 2023a) Aligned via Q-Formers for video and audio Instruction-tuned multimodal model.
ChatVideo (Wang et al., 2023a) arXiv 2023 e.g., OmniVL(Wang et al., 2022a), InternVideo(Wang et al., 2022b) Text: ChatGPT(Wu et al., 2023b), Audio: e.g., Whisper(Radford et al., 2021) Tracklet-centric with ChatGPT reasoning Chat-based video understanding system.
VideoChat (Li et al., 2023a) arXiv 2023 EVA-CLIP ViT-G/14 (Sun et al., 2023) Text: StableVicuna(contributors, 2023), Audio: Whisper(Radford et al., 2021) Q-Former bridges visual features to LLMs for reasoning Chat-centric model.
VideoLLM (Chen et al., 2023b) arXiv 2023 e.g., I3D(Carreira and Zisserman, 2017), SlowFast (Feichtenhofer et al., 2019) Text: e.g., BERT(Devlin et al., 2019), T5(Raffel et al., 2020) Semantic translator aligns visual and text encodings Video sequence modeling using LLMs.
VAST (Chen et al., 2023a) NeurIPS 2023 EVA-CLIP ViT-G/14(Sun et al., 2023) Text: BERT(Devlin et al., 2019), Audio: BEATs(Chen et al., 2022b) Cross-attention layers Omni-modality foundational model.
Video-ChatGPT (Maaz et al., 2023) ACL 2023 CLIP ViT-L/14 (Radford et al., 2021) Text: Vicuna-v1.1(Liu et al., 2023) Spatiotemporal features projected via linear layer Integration of vision and language for video understanding.
Valley (Luo et al., 2023) arXiv 2023 CLIP ViT-L/14 (Radford et al., 2021) Text: StableVicuna(contributors, 2023) Projection layer LLM for video assistant tasks.
Macaw-LLM (Lyu et al., 2023) arXiv 2023 CLIP ViT-B/16 (Radford et al., 2021) Text: LLAMA-7B(Touvron et al., 2023), Audio: Whisper(Radford et al., 2021) Alignment module unifies multi-modal representations Multimodal integration using image, audio, and video inputs.
Auto-AD II (Han et al., 2023) CVPR 2023 CLIP ViT-B/32 (Radford et al., 2021) Text: BERT(Devlin et al., 2019) Cross-attention layers Movie description using vision and language.
Video-LLaVA (Lin et al., 2023) arXiv 2023 LanguageBind-Video (Zhu et al., 2023) Text: Vicuna v1.5(Chiang et al., 2023) MLP projection layer Unified visual representation learning for video.
GPT4Video (Wang et al., 2024l) ACMMM 2023 CLIP ViT-L/14 (Radford et al., 2021) Text: LLaMA 2(Touvron et al., 2023) Transformer-based cross-attention layer Video understanding with LLM-based reasoning.
LLaMA-VID (Li et al., 2025) ECCV 2023 CLIP ViT-L/14 (Radford et al., 2021) Text: Vicuna(Chiang et al., 2023) Context attention and linear projector LLaMA-VID for visual-textual alignment in video.
InternVideo2 (Wang et al., 2024h) ECCV 2023 InternVL-6B(Chen et al., 2024g), VideoMAE V2 (Wang et al., 2023b) Text: BERT-Large(Devlin et al., 2019), Audio: BEATs(Chen et al., 2022b) Q-Former aligns multi-modal embeddings Foundation model for multimodal video understanding.
COSMO (Wang et al., 2024i) arXiv 2024 CLIP ViT-L/14(Radford et al., 2021) Text: OPT-IML(Iyer et al., 2022)/RedPajama(Together.xyz, 2023)Mistral(Jiang et al., 2023) Gated cross-attention Contrastive-streamlined multimodal model.
VTimeLLM (Huang et al., 2024) CVPR 2024 CLIP ViT-L/14 (Radford et al., 2021) Text: Vicuna(Chiang et al., 2023) Linear layer Temporal video understanding enhanced with LLMs.
VILA (Lin et al., 2024) CVPR 2024 CLIP ViT-L/336px(Radford et al., 2021) Text: LLaMA-2-7B/13B(Touvron et al., 2023) Linear layer Vision-language model.
Video ReCap (Islam et al., 2024) CVPR 2024 TimeSformer (Bertasius et al., 2021) Text: GPT-2(Radford et al., 2019) Cross-attention layers Recursive hierarchical captioning model
OmniViD (Wang et al., 2024c) CVPR 2024 VideoSwin (Liu et al., 2022b) Text: BART(Lewis et al., 2019) MQ-Former Generative model for universal video understanding.
VTG-LLM (Guo et al., 2024a) arXiv 2024 EVA-CLIP ViT-G/14(Sun et al., 2023) Text: LLaMA-2-7B(Touvron et al., 2023) Projection layer Enhanced video temporal grounding.
AutoAD III (Han et al., 2024) CVPR 2024 EVA-CLIP ViT(Sun et al., 2023) Text: GPT-3.5-turbo Shared Q-Former Video description enhancement with LLMs.
LAVAD (Zanella et al., 2024) CVPR 2024 BLIP-2 ViT-L/14, ImageBind (Girdhar et al., 2023a) Text: Llama-2-13b-chat(Touvron et al., 2023) Converts video features into textual prompts for LLMs Training-free video anomaly detection using LLMs.
MA-LMM (He et al., 2024) CVPR 2024 EVA-CLIP ViT-G/14 Text: Vicuna(Chiang et al., 2023) A trainable Q-Former Memory-augmented large multimodal model.
MiniGPT4-Video (Ataallah et al., 2024) arXiv 2024 EVA-CLIP ViT(Sun et al., 2023) Text: LLaMA 2(Touvron et al., 2023) Concatenates visual tokens and projects into LLM space Video understanding with visual-textual token interleaving.
PLLaVA (Xu et al., 2024) arXiv 2024 CLIP ViT-L/14 (Radford et al., 2021) Text: LLAMA-7B(Touvron et al., 2023) MM projector with adaptive pooling Parameter-free extension for video captioning tasks.
V2Xum-LLaMA (Hua et al., 2024) arXiv 2024 CLIP ViT-L/14 (Radford et al., 2021) Text: LLaMA 2(Touvron et al., 2023) Vision adapter Video summarization using temporal prompt tuning.
VideoChat2 (Li et al., 2024) CVPR 2024 UMT-L(Liu et al., 2022a) Text: Vicuna(Chiang et al., 2023) Linear projection A comprehensive multi-modal video understanding benchmark.
MotionLLM (Chen et al., 2024a) arXiv 2024 LanguageBind(Zhu et al., 2023), VQ-VAE(Zhang et al., 2023b) Text: Vicuna(Chiang et al., 2023) Modality translator: Motion / Video translator Understanding human behaviors from human motions and videos.
VideoGPT+ (Maaz et al., 2024) arXiv 2024 CLIP ViT-L/14(Radford et al., 2021), InternVideo2(Wang et al., 2024h) Text: Phi-3-Mini-3.8B(Abdin et al., 2024) MLP Enhanced video understanding.
EmoLLM (Yang et al., 2024c) arXiv 2024 CLIP ViT-L/14(Radford et al., 2021) Text: Vicuna-v1.5(Chiang et al., 2023), Audio: Whisper(Radford et al., 2021) Multi-perspective visual projection Multimodal emotional understanding with improved reasoning.
Holmes-VAD (Zhang et al., 2024c) arXiv 2024 LanguageBind ViT-L/14(Zhu et al., 2023) Text: LLaMA3-Instruct-70B(AI, 2024) Temporal sampler Multimodal LLM for video anomaly detection.
ShareGPT4Video (Chen et al., 2024f) arXiv 2024 CLIP ViT-L/14(Radford et al., 2021) Text: Mistral-7B-Instruct-v0.2(Jiang et al., 2023) MLP Precise and detailed video captions with hierarchical prompts.
Vriptor (Yang et al., 2024a) arXiv 2024 EVA CLIP ViT-L/14(Sun et al., 2023) Text: ST-LLM(Liu et al., 2025), Audio: Whisper(Radford et al., 2021) Scene-level sequential alignment Vriptor for dense video captioning.
VideoLLaMA 2 (Cheng et al., 2024) arXiv 2024 CLIP ViT-L/14(Radford et al., 2021) Text: LLAMA 1.5(Liu et al., 2024b), Audio: BEATs(Chen et al., 2022b) Spatial-Temporal Convolution (STC) connector Advancing spatial-temporal modeling and audio understanding.
VideoLLM-online (Chen et al., 2024b) CVPR 2024 CLIP ViT-L/14(Radford et al., 2021) Text: Llama-2-Chat(Touvron et al., 2023)/Llama-3-Instruct(AI, 2024) MLP projector Online video large language model for streaming video.
Video-CCAM (Fei et al., 2024) arXiv 2024 SigLIP-SO400M(Zhai et al., 2023) Text: Phi-3-4k-instruct(Abdin et al., 2024)/ Yi-1.5-9B-Chat(AI et al., 2024) Cross-attention-based projector Causal cross-attention masks for short and long videos.
LongVA (Zhang et al., 2024a) arXiv 2024 CLIP ViT-L/336px (Radford et al., 2021) Text: Qwen2-Extended(Bai et al., 2023; Yang et al., 2024b) MLP Long context video understanding.
InternLM-XComposer-2.5(Zhang et al., 2024b) arXiv 2024 CLIP ViT-L/14 (Radford et al., 2021) Text: InternLM2-7B(Cai et al., 2024), Audio: Whisper(Radford et al., 2021) MLP Long-context LVLM supporting ultra-high-resolution video tasks.
VITA (Fu et al., 2024b) arXiv 2024 InternViT-300M (Chen et al., 2024c; Gao et al., 2024; Chen et al., 2024e, g) Text: Mixtral-8x7B (Jiang et al., 2024), Audio: Mel Filter Bank MLP Open-source interactive multimodal LLM.
Kangaroo (Liu et al., 2024d) arXiv 2024 EVA-CLIP-L (Sun et al., 2023) Text: Llama-3-8B-Instruct(AI, 2024) Multi-modal projector Video-language model supporting long-context video input.
Qwen2-VL (Wang et al., 2024a) arXiv 2024 CLIP ViT-L/14(Radford et al., 2021) Text: Qwen2-7B(Bai et al., 2023; Yang et al., 2024b) Cross-attention modules Vision-language model for multimodal tasks.
Oryx (Liu et al., 2024a) arXiv 2024 OryxViT(Liu et al., 2024a; Zhai et al., 2023) Text: Qwen2-7B/32B(Bai et al., 2023; Yang et al., 2024b) Cross-attention Spatial-temporal model for high-resolution understanding.
Video-XL (Shu et al., 2024) arXiv 2024 CLIP ViT-L(Radford et al., 2021) Text: Qwen2-7B(Bai et al., 2023; Yang et al., 2024b) Visual-language projector Long-context video understanding model.
SlowFocus (Nie et al., 2024) NeurIPS 2024 CLIP ViT-L/14(Radford et al., 2021) Text: Vicuna-7B v1.5(Zheng et al., 2023) Visual adapter (projector layer) Fine-grained temporal understanding in video LLM.
VideoStudio (Long et al., 2024) ECCV 2024 CLIP ViT-H/14 (Radford et al., 2021) Text: CLIP ViT-H/14(Radford et al., 2021) Cross-attention modules Multi-scene video generation.
VideoINSTA (Liao et al., 2024) arXiv 2024 CLIP ViT-L/14 (Radford et al., 2021) Text: Llama-3-8B-Instruct(AI, 2024) Self-reflective spatial-temporal fusion Zero-shot long video understanding model.
Loong (Wang et al., 2024m) arXiv 2024 Causal 3D CNN(Yu et al., 2024) Text: Standard text tokenizer Decoder-only autoregressive LLM with causal attention Autoregressive language models.
TRACE (Guo et al., 2024b) arXiv 2024 CLIP ViT-L(Radford et al., 2021) Text: Mistral-7B(Jiang et al., 2023) Task-interleaved sequence modeling & Adaptive head-switching Video temporal grounding via causal event modeling.
Apollo(Zohar et al., 2024) arXiv 2024 SigLIP-SO400M(Zhai et al., 2023), InternVideo2(Wang et al., 2024h) Text: Qwen2.5-7B(Bai et al., 2023; Yang et al., 2024b) Perceiver Resampler & Token Integration with Timestamps Video understanding model.
Table 1. Summary of latest multimodal video-LLMs and their interaction / fusion mechanisms.
By addressing these aspects, this paper seeks to advance our understanding of temporal reasoning in LLMs and pave the way for more robust, generalizable, and scalable solutions in video analysis. The insights offered here aim to engage researchers and practitioners alike, highlighting the importance of bridging the gap between static representations and dynamic reasoning in AI systems.
2. Related Work
The application of LLMs to video processing has attracted significant attention, owing to their capacity to bridge visual and textual modalities (Wang et al., 2024b). In this section, we review related work across four key areas: LLMs for video understanding, pretrained visual encoders, datasets for video understanding, and temporal reasoning in AI systems. We also highlight the distinct contributions of our work compared to prior studies.
Type Visual encoder Pretrained dataset
Image Normalizer-Free ResNet(Brock et al., 2021) ImageNet-1K(Deng et al., 2009)
CLIP ViT(Radford et al., 2021) WebImageText(Srinivasan et al., 2021)
EVA-CLIP ViT(Sun et al., 2023) LAION-2B(Schuhmann et al., 2022), COYO-700M(Byeon et al., 2022)
BLIP-2 ViT(Li et al., 2023b) WebImageText(Srinivasan et al., 2021)
SigLIP (e.g., ViT)(Zhai et al., 2023) WebLI(Chen et al., 2022a)
OryxViT(Liu et al., 2024a; Zhai et al., 2023) WebLI(Chen et al., 2022a)
Video TimeSformer(Bertasius et al., 2021) Kinetics-400(Kay et al., 2017), Kinetics-600(Carreira et al., 2018)
I3D(Carreira and Zisserman, 2017) HMDB51(Kuehne et al., 2011a), UCF101(Soomro, 2012), Kinetics-400(Kay et al., 2017)
SlowFast(Feichtenhofer et al., 2019) Kinetics-400(Kay et al., 2017), Kinetics-600(Carreira et al., 2018), Charades(Sigurdsson et al., 2016)
VideoSwin(Liu et al., 2022b) ImageNet-21K(Ridnik et al., 2021)
UMT(Liu et al., 2022a) Kinetics-400(Kay et al., 2017), AudioSet(Gemmeke et al., 2017)
LanguageBind(Zhu et al., 2023) VIDAL-10M(Zhu et al., 2023)
VideoMAE V2(Wang et al., 2023b) Kinetics-400(Kay et al., 2017), -600(Carreira et al., 2018), -700(Carreira et al., 2019), Something-Something V2(Goyal et al., 2017), AVA(Gu et al., 2018)
InternVL (e.g., InternViT-6B) (Chen et al., 2024e) Hybrid image-text datasets (e.g., LAION-en)(Cai et al., 2022; Changpinyo et al., 2021; Gu et al., 2022; Schuhmann et al., 2022; Sharma et al., 2018)
InternVideo2(Wang et al., 2024h) Hybrid video datasets (e.g., Kinetics-400(Kay et al., 2017), InternVid(Wang et al., 2024e))
Table 2. Visual encoders with their pretrained datasets.
LLMs for video understanding. LLMs have demonstrated remarkable versatility in video-related tasks by incorporating multimodal learning frameworks (Wu et al., 2023a). Notable works, such as Flamingo (Alayrac et al., 2022) by DeepMind, integrate visual and textual modalities for tasks like video captioning and video question answering (QA). Flamingo uses a cross-modal attention mechanism to align spatiotemporal video embeddings with text representations, showcasing the potential of LLMs in multimodal fusion. Other models (Yuan et al., 2021; Girdhar et al., 2022; Wang et al., 2022a; Girdhar et al., 2023b), including OmniVL (Wang et al., 2022a) and Florence (Yuan et al., 2021), explore unified architectures that handle images, videos, and text simultaneously, reducing reliance on domain-specific encoders (Wang et al., 2019c; Wang and Koniusz, 2021; Wang et al., 2021a; Koniusz et al., 2021; Wang and Koniusz, 2022b, a, 2023, 2024; Wang et al., 2024k, j). However, these works primarily focus on improving task performance without a deep analysis of how LLMs handle temporal dynamics, leaving their capacity for explicit time reasoning largely unexamined.
While prior studies (Dhingra et al., 2022; Jain et al., 2023; Bagad et al., 2023; Tan et al., 2024; Wang et al., 2024f; Gurnee and Tegmark, 2024) primarily emphasize task performance, we specifically investigate whether LLMs truly understand the concept of time. Our work explores the interplay between spatiotemporal embeddings and LLM frameworks, providing a deeper analysis of their temporal reasoning capabilities.
Pretrained visual encoders in multimodal learning. The success of LLMs in video understanding often hinges on the use of pretrained visual encoders (Lin et al., 2024; Chen et al., 2024c; Wang et al., 2024d; Gao et al., 2024). Models like CLIP (Radford et al., 2021), ResNet (He et al., 2016), and Vision Transformers (ViT) (Dosovitskiy et al., 2021) are frequently used for spatial feature extraction, while video-specific encoders such as I3D (Carreira and Zisserman, 2017), SlowFast (Feichtenhofer et al., 2019), TimeSformer (Bertasius et al., 2021), and Video Swin Transformer (Liu et al., 2022b) extract spatiotemporal features. These encoders are trained on large-scale datasets like ImageNet (Deng et al., 2009; Ridnik et al., 2021) and Kinetics (Kay et al., 2017; Carreira et al., 2018, 2019), enabling them to capture fine-grained features for downstream tasks. Table 2 presents an overview of widely used pretrained visual encoders along with their corresponding training datasets. Figure 2 compares the performance of visual encoders, showcasing image encoders evaluated on ImageNet-1K(Deng et al., 2009) and video encoders assessed on Kinetics-400 and Something-Something V2(Goyal et al., 2017). While the modularity of these encoders facilitates efficient system design, their reliance on general-purpose pretrained features poses limitations, particularly in domain-specific tasks and long-term temporal reasoning (Wang et al., 2019c; Wang and Koniusz, 2021; Koniusz et al., 2021; Qin et al., 2022; Wang and Koniusz, 2024; Wang et al., 2024k). Existing works often treat encoders as static components, overlooking the potential benefits of jointly optimizing encoders and LLMs for temporal understanding (Shi et al., 2024; Wang et al., 2023a).
Unlike works that use pretrained encoders as black-box components (Shu et al., 2023), we examine their limitations, including their bias toward short-term dependencies and their challenges in generalizing to abstract temporal concepts. We propose pathways for jointly optimizing encoders and LLMs to address these issues.
(a) Image encoders.
(b) Video encoders.
Figure 2. Performance comparison of visual encoders. (Left): Image classification accuracy for various image encoders pretrained and fine-tuned on the ImageNet-1K dataset. (Right): Action recognition accuracy for different video encoders pretrained and fine-tuned on the Kinetics-400 and Something-Something V2 datasets.
Datasets for video understanding. Datasets are fundamental to advancing video understanding, particularly for assessing the temporal reasoning abilities of LLMs (Girdhar and Ramanan, 2019; Zhou et al., 2018; Jang et al., 2019). Current datasets often focus on action recognition, video captioning, or question-answering, capturing spatiotemporal patterns and semantic connections (Goyal et al., 2017; Jang et al., 2017; Rohrbach et al., 2017). However, many datasets emphasize short-term motions or provide only surface-level annotations, lacking temporal details such as event order, causality, or duration (Kuehne et al., 2011b; Soomro, 2012; Kay et al., 2017; Wang et al., 2019d).
While multimodal datasets, combining video with text or audio, offer opportunities for LLMs to align spatiotemporal and semantic reasoning, challenges remain (Wang et al., 2024g; Fu et al., 2024a). These include limited diversity in scenarios, imprecise annotations, and the difficulty of representing long-term dependencies (Wang, 2023; Zhu et al., 2024). To truly evaluate and enhance LLMs’ temporal understanding, future datasets must provide richer, more diverse temporal annotations and robust benchmarks across varied domains (Nie et al., 2024). Table 3 provides a comprehensive summary of video datasets used for tasks such as action recognition, video question answering (QA), video captioning, video retrieval, and anomaly detection.
In this work, we emphasize the role of datasets in shaping temporal understanding. We analyze the shortcomings of current video datasets, including their lack of temporal annotations, short-term bias, and limited diversity, and propose directions for improving dataset design.
Task Dataset Year Source # Videos Modality Avg. length (s) Temporal annotation Description
Action Recognition HMDB51(Kuehne et al., 2011b) 2011 YouTube 6,766 Video 3~4 No Daily human actions
UCF101(Soomro, 2012) 2012 YouTube 13,320 Video+Audio 7.21 No Human actions (e.g., sports, daily activities)
ActivityNet(Caba Heilbron et al., 2015) 2015 YouTube 27,801 Video+Text 300~1200 Temporal extent provided Human-centric activities
Charades(Sigurdsson et al., 2016) 2016 Crowdsourced 9,848 Video+Text 30.1 Start and end timestamps provided Household activities
Kinetics-400(Kay et al., 2017) 2017 YouTube 306,245 Video 10 No Human actions (e.g., sports, tasks)
AVA(Gu et al., 2018) 2018 Movies 430 Video Variable Start and end timestamps provided Action localization in movie scenes
Kinetics-600(Carreira et al., 2018) 2018 YouTube 392,622 Video 10 No Human actions (e.g., sports, tasks)
Something-Something V2(Goyal et al., 2017) 2018 Crowdsourced 220,847 Video 2~6 Weak Human-object interactions
EPIC-KITCHENS(Damen et al., 2018) 2018 Participant kitchens 432 Video+Text+Audio ~458 Start and end timestamps provided Large-scale egocentric cooking dataset
COIN(Tang et al., 2019) 2019 YouTube 11,827 Video+Text 141.6 Start and end timestamps provided Comprehensive instructional tasks (e.g., cooking, repair)
Kinetics-700(Carreira et al., 2019) 2019 YouTube 650,317 Video 10 No Expanded version of Kinetics-400 and Kinetics-600
EPIC-KITCHENS-100(Damen et al., 2022) 2020 Participant kitchens 700 Video+Text+Audio ~514 Start and end timestamps provided Large-scale egocentric cooking dataset
Ego4D(Grauman et al., 2022) 2021 Wearable Cameras 3,850 hours Video+Text+Audio Variable Start and end timestamps provided First-person activities and interactions
VidSitu(Sadhu et al., 2021) 2021 YouTube 29,000 Video+Text ~10 Temporal extent for events provided Event-centric and causal activity annotations
Video QA MovieQA(Tapaswi et al., 2016) 2016 Multiple platforms 408 Video+Text 202.7 Start and end timestamps provided QA for movie scenes
TGIF-QA(Jang et al., 2017) 2016 Tumblr GIFs 56,720 Video+Text 3~5 Action timestamps provided QA over social media GIFs
MSVD-QA(Xu et al., [n. d.]) 2017 YouTube 1,970 Video+Text 27.5 Start and end timestamps provided QA for actions description
MSRVTT-QA(Xu et al., [n. d.]) 2017 YouTube 10,000 Video+Text 15~30 Weak QA across diverse scenes
TVQA(Lei et al., 2018) 2019 TV Shows 21,793 Video+Text 60~90 Start and end timestamps provided QA over medical dramas, sitcoms, crime shows
ActivityNet-QA(Yu et al., 2019) 2019 YouTube 5,800 Video+Text 180 Implicit (derived from ActivityNet) QA for human-annotated videos
How2QA(Li et al., 2020) 2020 HowTo100M (YouTube) 22,000 Video+Text 60 Temporal extent provided QA over instructional videos
YouCookQA(Wang et al., 2021b) 2021 YouCook2 (YouTube) 2,000 Video+Text 316.2 Temporal boundaries provided Cooking-related instructional QA
STAR(Wu et al., 2024) 2021 Human activity datasets 22,000 Video+Text Variable Action-level boundaries provided QA over human-object interactions
MVBench(Li et al., 2024) 2023 Public datasets 3,641 Video+Text 5~35 Start and end timestamps provided Multi-domain QA (e.g., sports, indoor scenes)
EgoSchema(Mangalam et al., 2023) 2023 Ego4D (Wearable Cameras) 5,063 Video+Text 180 Timestamped narrations provided Long-form egocentric activities
Video Captioning YouCook(Das et al., 2013) 2013 YouTube 88 Video+Text 180~300 Weak Cooking instructional videos
MSR-VTT(Xu et al., 2016) 2016 YouTube 7,180 Video+Text+Audio 10~30 Weak General scenarios (e.g., sports, transport)
ActivityNet Captions(Krishna et al., 2017) 2017 YouTube 20,000 Video+Text 180 Start and end timestamps provided Dense captions for human-centered activities
VATEX(Wang et al., 2019d) 2019 YouTube 41,250 Video+Text ~10 Weak Multilingual descriptions with English-Chinese parallel captions
HowTo100M(Miech et al., 2019) 2019 YouTube 1.22M Video+Text+Audio 390 Subtitle timestamps provided Instructional video captions
TVC(Lei et al., 2020) 2020 TV Shows 108,965 Video+Text 76.2 Start and end timestamps provided Multimodal video captioning dataset
Video Retrieval LSMDC(Rohrbach et al., 2017) 2015 Movies 118,114 Video+Text 4.8 Start and end timestamps provided Large-scale dataset for movie description tasks
DiDeMo(Hendricks et al., 2017) 2017 Flickr (YFCC100M) 10,464 Video+Text 27.5 Start and end timestamps provided Moment localization in diverse, unedited personal videos
FIVR-200K(Kordopatis-Zilos et al., 2019) 2019 YouTube 225,960 Video ~120 Start and end timestamps provided Large-scale incident video retrieval dataset with diverse news events
TVR(Lei et al., 2020) 2020 TV Shows 21,793 Video+Text 76.2 Start and end timestamps provided Video-subtitle multimodal moment retrieval dataset
TextVR(Wu et al., 2023c) 2023 YouTube 10,500 Video+Text 15 Weak Cross-modal video retrieval with text reading comprehension
EgoCVR(Hummel et al., 2024) 2024 Ego4D 2,295 Video+Text 3.9~8.1 Weak Egocentric dataset for fine-grained composed video retrieval
Anomaly Detection Subway Entrance(Adam et al., 2008) 2008 Surveillance cameras 1 Video 4,800 No Crowd monitoring for unusual event detection at subway entrances
Subway Exit(Adam et al., 2008) 2008 Surveillance cameras 1 Video 5,400 No Crowd monitoring for unusual event detection at subway exits
CUHK Avenue(Lu et al., 2013) 2013 Surveillance cameras 15 Video 120 No Urban avenue scenes with anomalies like running, loitering, etc.
Street Scene(Ramachandra and Jones, 2020) 2020 Urban street surveillance 81 Video 582 Spatial and temporal bounding boxes Urban street anomalies, e.g., jaywalking, loitering, illegal parking, etc.
XD-Violence(Wu et al., 2020) 2020 Movies and in-the-wild scenes 4,754 Video+Audio ~180 Start and end timestamps provided Multimodal violence detection covering six violence types
CUVA(Du et al., 2024) 2024 YouTube, Bilibili 1,000 Video+Text ~117 Start and end timestamps provided Causation-focused anomaly understanding across 42 anomaly types
MSAD(Zhu et al., 2024) 2024 Online Surveillance 720 Video ~20 Frame-level annotations in test set Multi-scenario dataset with 14 scenarios
Multimodal video tasks VIDAL-10M(Zhu et al., 2023) 2023 Multiple platforms 10M Video+Infrared+Depth+Audio+Text ~20 Weak Multi-domain retrieval dataset
Video-MME(Fu et al., 2024a) 2024 YouTube 900 Video+Text+Audio 1017.9 Temporal ranges via certificate length Comprehensive evaluation benchmark across many domains
Table 3. Comprehensive overview of video datasets across tasks.
Temporal reasoning in video AI. Temporal reasoning is critical for tasks such as action recognition, video summarization, and temporal event ordering (Wang, 2021; Wang et al., 2019b, c; Wang and Koniusz, 2021; Wang et al., 2021a; Koniusz et al., 2021; Wang and Koniusz, 2022b, a; Qin et al., 2022; Wang and Koniusz, 2023, 2024; Wang et al., 2024k, j; Raj et al., 2024; Chen et al., [n. d.], 2024d, 2024h; Zhu et al., 2024; Ding et al., 2024). Classical approaches rely on recurrent neural networks (RNNs) (Sherstinsky, 2020), 3D convolutional networks (3D-CNNs) (Tran et al., 2015; Carreira and Zisserman, 2017), and attention mechanisms (Vaswani et al., 2017; Chen et al., [n. d.]; Wang and Koniusz, 2023) to model temporal dependencies. More recently, transformers and temporal tokenization techniques have been explored for long-term video understanding (Bertasius et al., 2021; Patrick et al., 2021). Despite these advances, the explicit modeling of abstract temporal concepts such as causality, sequence progression, and event duration remains underexplored. LLM-based approaches typically use spatiotemporal embeddings from pretrained encoders but fall short in demonstrating robust temporal reasoning capabilities, especially for complex, real-world video scenarios (Wang, 2023; Zhu et al., 2024).
While many studies report empirical results (Dhingra et al., 2022; Jain et al., 2023; Bagad et al., 2023; Tan et al., 2024; Wang et al., 2024f; Gurnee and Tegmark, 2024; Liu et al., 2024c), we go further by offering actionable insights for advancing LLM-based temporal reasoning. These include joint training approaches, better dataset curation, and innovative multimodal fusion techniques. By addressing these gaps, our work not only advances the understanding of temporal reasoning in LLMs but also provides a roadmap for future research, distinguishing it from prior efforts(Tang et al., 2023; Nguyen et al., 2024; Zhou et al., 2024) in this rapidly evolving field.
3. Analysis and Discussion
Can LLMs understand the concept of time? LLMs exhibit several strengths in processing temporal information. Trained on textual data containing narrations or instructions, they can infer temporal relationships through contextual cues such as “first”, “then”, and “after”. When paired with video encoders, LLMs can process spatiotemporal embeddings, enabling tasks like action recognition and temporal event ordering.
However, LLMs lack direct temporal awareness. Standard models do not inherently model the flow of time unless explicitly trained on sequential video data. Instead, they rely on external encoders to provide temporal structure. Capturing long-term dependencies over extended video sequences is another challenge, as LLMs often operate on tokenized inputs within limited context windows. Additionally, video encoders like 3D CNNs (Tran et al., 2015; Carreira and Zisserman, 2017) or video transformers (Patrick et al., 2021; Wang et al., 2023b), which act as the “eyes” of the system, may excel in capturing motion patterns but struggle to generalize abstract temporal concepts like causality or duration.
A significant limitation lies in the representation of visual time. Unlike textual representations, visual cues require explicit modeling of motion and event transitions. This ambiguity underscores the need for improved temporal modeling in both encoders and LLMs.
LLMs applied to videos using pretrained visual encoders. Most existing approaches use pretrained image or video encoders to extract visual or spatiotemporal information, rather than designing entirely new encoders (Table 1 and 2). Pretrained image encoders such as CLIP (Radford et al., 2021), ResNet (He et al., 2016), and ViT (Vaswani et al., 2017) excel in capturing spatial information, while video encoders like I3D (Carreira and Zisserman, 2017), SlowFast (Feichtenhofer et al., 2019), TimeSformer (Bertasius et al., 2021), and Video Swin Transformer (Liu et al., 2022b) are widely used for spatiotemporal feature extraction. These encoders, trained on large-scale datasets such as ImageNet (Deng et al., 2009; Ridnik et al., 2021) or Kinetics (Kay et al., 2017; Carreira and Zisserman, 2017; Carreira et al., 2019), are adept at learning rich feature representations, which can then be fine-tuned for specific tasks (Wang et al., 2019c; Wang and Koniusz, 2021; Koniusz et al., 2021; Wang, 2023; Wang and Koniusz, 2024; Wang et al., 2024k; Chen et al., [n. d.]; Ding et al., 2024; Zhu et al., 2024; Wang et al., [n. d.]; Chen et al., 2024d). Notably, video encoders incorporate mechanisms to model temporal dependencies, such as optical flow (Wang and Koniusz, 2024), Taylor videos (Wang et al., [n. d.]), and motion tracking (Chen et al., [n. d.]; Raj et al., 2024), addressing a challenge that LLMs alone cannot handle effectively. Figure 2 illustrates the performance of popular visual encoders on two key tasks: image classification (ImageNet-1K) and video action recognition (Kinetics-400 and Something-Something V2).
The use of pretrained visual encoders offers practical advantages. These encoders are optimized for handling visual and spatiotemporal features, reducing computational overhead and enabling faster convergence. Their modular design also allows them to function as “plug-and-play” components, seamlessly integrating with LLMs for multimodal learning. This modularity ensures that LLM-based frameworks remain adaptable and scalable across diverse applications.
However, pretrained encoders are not without limitations. First, encoders trained on general datasets may underperform on domain-specific video tasks (Zhu et al., 2024; Chen et al., 2024d). Second, many pretrained encoders prioritize spatial over temporal information, necessitating additional modules, such as temporal transformers, to capture complex temporal dynamics (Chen et al., 2024a; Zhang et al., 2024c). Lastly, large-scale video datasets required for training such encoders are costly to annotate, limiting their ability to encompass diverse or abstract video content.
Some recent efforts, such as DeepMind’s Flamingo (Alayrac et al., 2022) and unified architectures like Florence (Yuan et al., 2021) and OmniVL (Wang et al., 2022a), explore custom encoders optimized for multimodal learning. These models aim to balance performance across multiple modalities (image, video, and text) without relying heavily on separate pretrained components (Wang et al., 2019c; Wang and Koniusz, 2021; Wang, 2023).
How encoders and LLMs interact? Encoders play a crucial role in preprocessing video frames or sequences to extract visual or spatiotemporal features, which are then transformed into a format compatible with LLMs, often as token embeddings. For example, video transformers process sequences of video frames, while text encoders like CLIP encode textual inputs (Yang et al., 2023; Zhang et al., 2023a; Ataallah et al., 2024). These features are projected into a shared representation space to align different modalities, such as visual and textual data.
(a) Interaction/fusion.
(b) Modalities.
Figure 3. The distributions of interaction/fusion mechanisms and data modalities in $6$ closely related video-LLMs from January 2024 to December 2024. (Left): Fusion mechanisms are classified into five categories: Cross-attention (e.g., cross-attention modules, gated cross-attention), Projection layers (e.g., linear projection, MLP projection), Q-Former-based methods (e.g., Q-Former aligns multi-modal embeddings, Trainable Q-Former), Motion/Temporal-Specific mechanisms (e.g., temporal samplers, scene-level sequential alignment), and Other Methods (e.g., Tracklet-centric, Perceiver Resampler, MQ-Former). (Right): The distribution of data modalities used in these video-LLMs, with text modalities appearing across all models. Note that a model may use multiple fusion methods and/or data modalities.
Fusion mechanisms are pivotal in enabling LLMs to process multimodal inputs. Encoded features are tokenized and treated as input tokens for the LLM. Attention mechanisms, such as cross-modal attention (Zhao et al., 2023; Chen et al., 2023a), are then used to integrate and relate features from different encoders. Positional embeddings encode spatial and temporal positions of video features, helping the LLM model sequence dependencies. For instance, cross-modal transformers align visual and textual representations, while LLMs refine these embeddings to generate outputs like action labels, video descriptions, or anomaly detection results (Fei et al., 2024).
Table 1 highlights the interaction/fusion mechanisms used in the latest video-LLMs under the ‘Interaction / Fusion mechanism’ column. Figure 3(a) visualizes the distribution of various interaction and fusion techniques, while Figure 3(b) showcases the modalities used in closely related works from 2024 (notably, text modalities are consistently used across these studies). As shown in these plots, the majority of works use projection layers (e.g., linear projection (Li et al., 2024), MLP projection (Zhang et al., 2024a, b; Fu et al., 2024b), semantic translators (Chen et al., 2023b, 2024a)) and cross-attention mechanisms (e.g., cross-modal attention (Yang et al., 2023), gated cross-attention (Wang et al., 2024i)) to facilitate interaction between encoders and LLMs. Temporal-specific mechanisms, such as temporal sampler (Zhang et al., 2024c) and scene-level sequential alignment (Yang et al., 2024a), are also used. Emerging video-LLMs are beginning to explore novel fusion mechanisms, including tracklet-centric approach (Wang et al., 2023a), Perceiver resamplers (Alayrac et al., 2022; Zohar et al., 2024), and MQ-Former (Wang et al., 2024c).
However, several challenges arise in this interaction. Encoders must output features in a format compatible with LLMs, requiring careful dimension alignment and embedding space mapping. Additionally, processing video data generates a substantial volume of information, which can strain the LLM’s capacity. Capturing long-term dependencies across extended video sequences also remains a challenge, even with support from pretrained encoders.
Bridging the gap between raw video data and temporal reasoning. The interaction between LLMs and encoders helps bridge the gap between raw video data and higher-level temporal reasoning. Pretrained encoders extract spatiotemporal embeddings that encapsulate low-level motion cues, such as velocity (Wang et al., [n. d.]) and trajectory (Patrick et al., 2021), as well as higher-level temporal patterns like scene transitions and sequence progression. These embeddings provide the foundation for LLMs to interpret complex temporal concepts such as causality, event progression, and anticipation.
LLMs use attention mechanisms to prioritize spatiotemporal features, aligning them with semantic or task-specific contexts. Encoders often supply frame-level features as temporal tokens, enabling the LLM to model dependencies and transitions across frames. However, significant challenges remain in achieving comprehensive temporal understanding. Encoders frequently focus on short-term motion patterns, neglecting long-term dependencies (Wang et al., 2021a; Wang and Koniusz, 2022a; Wang et al., 2024j). Similarly, datasets used for encoder pretraining often lack diversity and fail to represent abstract temporal relationships effectively (Wang, 2023). Table 3 provides a comparison of datasets based on average video length, data source, modalities, and the number of videos.
LLMs also face inherent limitations. Temporal embeddings require careful tokenization to preserve sequence information when input into LLMs (see Table 1). Additionally, LLMs pretrained on static datasets such as text or images may lack the dynamic reasoning capabilities required for video tasks (Han et al., 2023). Combining spatial, temporal, and semantic information without losing critical cues remains a complex challenge (Chen et al., 2024d, [n. d.]; Raj et al., 2024), further compounded by the computational expense of processing long video sequences while retaining both global and local temporal details.
Figure 4. Performance (accuracy) comparison of recent video-LLMs on the Video-MME benchmark.
Video datasets: an enabler or bottleneck? Video datasets are foundational to LLMs on video tasks, yet they often act as a bottleneck (Wang, 2023). Action recognition datasets, such as Kinetics (Kay et al., 2017; Carreira et al., 2018) and Something-Something V2 (Goyal et al., 2017), are effective for analyzing short-term motion patterns but lack the temporal annotations necessary for reasoning about action sequences or causal relationships. Compared to the Something-Something V2 dataset, the Kinetics datasets exhibit a stronger spatial bias, as demonstrated in (Patrick et al., 2021; Girdhar et al., 2022, 2023b). Similarly, video QA datasets like TVQA (Lei et al., 2018) and How2QA (Li et al., 2020) align well with LLM architectures but often rely on scripted scenarios that limit generalizability to real-world tasks.
Video captioning datasets like MSR-VTT (Xu et al., 2016) and ActivityNet Captions (Krishna et al., 2017) enable multimodal learning by fusing video and text embeddings. However, captions are often superficial and fail to probe deeper temporal reasoning. Long-term video understanding datasets, such as Ego4D (Grauman et al., 2022) and Charades (Sigurdsson et al., 2016), focus on extended activities and interactions, offering a richer testing ground for temporal reasoning. Nonetheless, LLMs often struggle with the scale and complexity of such datasets, given their limited context windows.
Multimodal datasets like HowTo100M (Miech et al., 2019) and COIN (Tang et al., 2019) align video content with auxiliary modalities, providing opportunities for pretraining in a multimodal setup. However, the inherent noise and lack of temporal annotations in these datasets can hinder performance. To advance temporal understanding in LLMs, datasets must evolve to include richer temporal annotations, long-term dependencies, and diverse real-world scenarios. The latest multimodal datasets include VIDAL-10M (Zhu et al., 2023) and Video-MME (Fu et al., 2024a).
Table 3 summarizes popular datasets used across various video tasks, including action recognition, anomaly detection, video question answering, captioning, and retrieval, as well as some recent multimodal video understanding datasets.
Figure 5. Performance comparison of recent video-LLMs on video QA benchmarks. Models using pretrained video encoders (e.g., Video-LLaVA and VideoChat2) are marked with squares, while models using pretrained image encoders are represented by circles.
State-of-the-Art video LLMs. Recent advancements in video LLMs have significantly enhanced the processing of spatiotemporal information (Chen et al., 2024b; Nie et al., 2024; Li et al., 2024). Traditional LLMs, focused on textual data, often struggle with temporal dynamics in video (Liu et al., 2024c). Models like VideoChat2 (Li et al., 2024) and SlowFocus (Nie et al., 2024) are pushing the boundaries by integrating temporal reasoning with video analysis. VideoChat2 enables real-time multimodal dialogue, processing video sequences to answer questions about actions, events, and causal relationships, making it highly effective for interactive applications. SlowFocus improves fine-grained temporal understanding, capturing long-term dependencies and transitions within video, essential for tasks like video summarization and anomaly detection. TimeSformer (Bertasius et al., 2021) further refines this by using attention mechanisms to simultaneously model spatial and temporal features, enhancing video understanding in complex scenarios like action recognition.
Additionally, models, e.g., VideoLLM-Online (Chen et al., 2024b) and VideoBERT (Sun et al., 2019), focus on continuous learning from video streams, allowing for real-time updates and better adaptability to dynamic content. These models are crucial for applications requiring ongoing video analysis, such as surveillance, event detection, and interactive media. Flamingo (Alayrac et al., 2022) takes multimodal learning a step further by combining visual, textual, and temporal data, offering a more holistic approach to video processing. ActionFormer (Zhang et al., 2022), on the other hand, specializes in action recognition through long-range temporal dependencies, making it effective for tasks like sports video analysis and human-computer interaction. These advancements reflect a significant leap in video LLM capabilities, making them better equipped for real-world video understanding, interaction, and analysis.
Figures 4, 5, and 6 present comparisons of recent popular video-LLMs across multiple video tasks, including multimodal video understanding, video QA, video retrieval, and video captioning. As shown in these figures, no single video-LLM excels across all tasks: (i) a comprehensive evaluation system covering all video tasks is lacking, and (ii) most video-LLMs are designed to address only a subset of these challenges.
Fair evaluation is needed. Evaluations and comparisons of video-LLMs are often conducted inconsistently (see Figures 4, 5, and 6), which can result in unfair assessments and misleading conclusions (Tang et al., 2023; Nguyen et al., 2024). A frequent issue is the comparison of video-LLMs, designed for multimodal reasoning, with traditional video models such as I3D (Carreira and Zisserman, 2017), SlowFast (Feichtenhofer et al., 2019), or Video Swin Transformer (Liu et al., 2022b), which are tailored for video-specific tasks like action recognition. While traditional models excel at spatiotemporal feature extraction due to their focused design, video-LLMs must simultaneously handle visual and linguistic alignment, which adds inherent complexity to their objectives. Directly comparing video-LLMs against such specialized models is therefore not entirely fair. Instead, video-LLMs should be systematically benchmarked against other video-LLMs or multimodal frameworks to better reflect their relative strengths and limitations in tasks like video action recognition, video captioning, or video QA.
Furthermore, inconsistencies in evaluation practices exacerbate the problem. Different models are often trained and tested on varying datasets, such as Kinetics-400 (Kay et al., 2017), Ego4D (Grauman et al., 2022), or HowTo100M (Miech et al., 2019), without standardized protocols for pretraining, finetuning, or testing. This creates biases in results and hampers fair comparisons. For instance, a video-LLM pretrained on massive multimodal datasets might show superior results simply due to larger data availability rather than architectural improvements. To address this, evaluations must adopt consistent and standardized benchmarks, training splits, and metrics. Frameworks that systematically assess multimodal alignment, temporal reasoning, and downstream task performance would help ensure transparency and comparability.
Finally, establishing fair evaluation practices requires prioritizing within-paradigm comparisons. For video action recognition, for example, models like VideoChat2 (Li et al., 2024) and VideoLLM-online (Chen et al., 2024b) should be compared against each other rather than with traditional video-only transformers. This approach highlights progress within the multimodal video understanding space and reveals areas for improvement, such as better temporal consistency or more efficient multimodal alignment. By addressing these challenges, fair and systematic evaluation will provide deeper insights into video-LLM capabilities and foster future advancements in the field.
Factors driving superior performance in video LLMs. The superior performance of certain video LLMs can be attributed to their ability to effectively integrate spatial, temporal, and semantic information, often through advanced architectures and training strategies. Models like SlowFocus (Nie et al., 2024) and VideoChat2 (Li et al., 2024) excel by incorporating fine-grained temporal reasoning and long-range dependencies, which are crucial for understanding complex video dynamics such as event progression and causal relationships. The use of hierarchical or multi-level attention mechanisms, seen in models like TimeSformer (Bertasius et al., 2021) and VideoLLM-Online (Chen et al., 2024b), enables them to capture both short-term motions and long-term narrative structures, addressing the temporal gaps that limit earlier models.
Additionally, models that use large-scale, multimodal pretraining on diverse video datasets, such as Flamingo, benefit from the cross-domain knowledge transfer between visual, textual, and temporal modalities, leading to a more holistic understanding of video content. These innovations enable the models to generalize better across different video tasks, including action recognition, video captioning, and dynamic scene interpretation.
(a) Video retrieval.
(b) Video captioning.
Figure 6. Performance comparison of recent video-LLMs on (a) video retrieval and (b) video captioning benchmarks.
4. Future Directions
Building on the preceding analysis and discussion, we outline below several promising future research directions for those interested in advancing video LLMs.
Overcoming dataset challenges for LLMs. Datasets remain a critical bottleneck in advancing LLM-based video systems. Addressing their limitations requires both creative solutions and resource investments:
  1. i.
    Temporal annotations and structure: Enriching datasets with temporal annotations, such as event order, duration, and causal relationships, is essential. Techniques like crowd-sourced annotations, AI-assisted labeling, or synthetic data generation could address the scarcity of such datasets.
  2. ii.
    Balancing scale and quality: While large-scale datasets like HowTo100M offer broad coverage, they often include noise and irrelevant data. Future efforts should focus on curating high-quality, balanced datasets that prioritize diversity and accuracy without sacrificing scale. Semi-supervised and unsupervised learning approaches could also mitigate the need for large annotated datasets.
  3. iii.
    Addressing short-term bias: The dominance of short video clips in existing datasets limits models’ ability to reason over extended sequences. Introducing datasets with long-term dependencies, such as episodic or procedural videos, would help train models to understand overarching narratives and transitions.
  4. iv.
    Expanding dataset diversity: Current datasets often focus on narrow domains, such as sports or cooking, which restricts model generalizability (Wang, 2023). Diverse datasets encompassing a wider range of cultural contexts, real-world scenarios, and tasks are essential to improve performance across applications. Collaborative initiatives between academic and industry stakeholders could accelerate the development of such datasets.
  5. v.
    Improving multimodal alignment: Misalignment between video frames and associated text or audio annotations introduces inconsistencies that can degrade learning quality. Future work could explore more precise alignment methods (Wang et al., 2021a; Wang and Koniusz, 2022a, b; Wang et al., 2024j), such as neural alignment models or reinforcement learning frameworks, to ensure that temporal and semantic signals are accurately correlated during training.
Enhancing temporal understanding. To improve temporal reasoning, joint training of encoders and LLMs is a promising direction. Models co-trained on datasets with temporal reasoning tasks can develop a deeper understanding of complex time-related concepts such as causality, event sequencing, and duration. Architectures like temporal transformers, recurrent neural networks, or hybrid systems that combine hierarchical and sequential processing should be further explored to handle both short-term dynamics and long-term dependencies in video data.
Explicit supervision for abstract temporal concepts through enriched annotations is another critical step (Ding et al., 2024). Annotated datasets with detailed temporal labels, covering relationships, transitions, and event causality, can significantly boost the temporal reasoning capacity of these systems. Furthermore, multimodal training on datasets that integrate video, text, audio, and temporal metadata could align spatiotemporal and semantic representations more effectively, enhancing the model’s ability to reason about time in real-world scenarios (Ding et al., 2024; Raj et al., 2024; Zhu et al., 2024).
Multimodal LLMs for holistic temporal reasoning. The development of truly multimodal LLMs capable of holistic temporal reasoning requires a synergistic interplay between spatiotemporal and semantic understanding. Future research should explore adaptive attention mechanisms that dynamically weigh temporal, spatial, and textual information based on context. Advances in memory-efficient architectures and progressive learning strategies could enable LLMs to process longer sequences without losing critical details.
Additionally, fine-tuning LLMs on domain-specific datasets or tasks, such as medical video analysis or video-based social interaction studies (Christophe et al., 2024; Jeong, 2024), could expand their applicability and temporal reasoning depth. Transfer learning approaches (Zhao et al., 2024), where models pretrained on general datasets are fine-tuned for specific temporal reasoning tasks, can also be effective in reducing resource demands.
Other modalities, such as depth videos (Wang, 2021; Wang et al., 2019a) and motion-specific data like skeletons (useful for analyzing human-related movements) (Wang et al., 2021a; Wang and Koniusz, 2022a, b, 2023; Wang et al., 2024j), optical flow (Wang and Koniusz, 2024), and Taylor videos (Wang et al., [n. d.]), can significantly enhance the performance of LLM frameworks in video processing tasks when their pretrained models are incorporated. Depth videos capture three-dimensional spatial information, offering a richer understanding of scene geometry. Skeleton data focuses on joint movements, making it particularly effective for applications like action recognition and gesture analysis. Optical flow and Taylor videos excel in capturing frame-to-frame motion changes, providing detailed temporal cues essential for understanding dynamic content.
By integrating these diverse modalities, LLMs can achieve a more comprehensive representation of motion dynamics and spatial structures, broadening their applicability to complex video-based challenges. Moreover, the inclusion of learned video motion prompts (Chen et al., [n. d.]), which highlight relevant movements within a scene, introduces a novel modality that further refines the system’s ability to process and interpret intricate video content.
Advancing visual encoders for multimodal learning. While most current LLM-based video systems use pretrained encoders for their efficiency and robust feature extraction, there is significant potential in designing novel encoders optimized specifically for multimodal learning. These encoders should aim to seamlessly integrate spatial, temporal, and semantic information into a unified framework, reducing the reliance on modular preprocessing steps. Future research could explore adaptive encoder architectures that dynamically adjust to varying video characteristics, such as scene complexity or temporal dynamics. Additionally, encoders tailored for specific domains, like healthcare, education, or autonomous driving, could enhance the accuracy and relevance of multimodal systems in specialized applications.
Ethical and practical considerations. As LLM-based video systems advance, addressing ethical concerns becomes increasingly important (Wang et al., [n. d.]; Zhu et al., 2024). Ensuring fairness and avoiding biases in datasets, particularly regarding cultural and contextual diversity, will be critical. Practical considerations like energy efficiency and the environmental impact of large-scale training should also guide future research. Lightweight models or distillation techniques could balance performance with computational sustainability.
Addressing these directions, researchers can unlock the full potential of LLMs in video-based applications, driving advancements in temporal reasoning, multimodal understanding, and real-world usability.
5. Conclusion
This work critically examines the temporal reasoning capabilities of large language models (LLMs) in video processing, identifying significant limitations in both models and datasets. While LLMs paired with pretrained visual encoders have achieved success in tasks such as action recognition, anomaly detection, and video summarization, they fall short in understanding long-term temporal dependencies. This stems from the encoders’ focus on short-term patterns, fragmented temporal cues, and challenges in aligning spatial, temporal, and semantic information. Additionally, existing datasets lack explicit temporal annotations, often focus on short clips over long sequences, and struggle with diversity and multimodal alignment, further hindering progress.
To unlock the full potential of LLMs in video processing, future research must address these gaps. This includes designing integrated frameworks to jointly train encoders and LLMs on temporal reasoning, enriching datasets with detailed annotations for long-term dependencies, and creating innovative architectures that fuse spatiotemporal and semantic information. By addressing these challenges, we can pave the way for systems that not only excel in video analysis but also advance broader applications requiring robust temporal comprehension.
Acknowledgements.
Xi Ding, a Research Assistant with the Temporal Intelligence and Motion Extraction (TIME) Lab at ANU, contributed to this work. This research was conducted under the supervision of Lei Wang.